# Litmus tests of the flat ${\Lambda\text{CDM}}$ model and model-independent measurement of $H_{0}r_{\text{d}}$ with LSST and DESI

### v1

@@ 3.2 Iterative Smoothing
The smoothing width is set to $\Delta=0.3$ following previous analyses in [41, 42, 39, 44, 43, 45, 46, 47].
We define the $\chi^{2}$ value of the reconstruction $\hat{y}_{n}(t)$ as
$\chi_{n}^{2}=\bm{\delta\mu}_{n}^{\intercal}\cdot\bm{C}_{\text{SN}}^{-1}\cdot\bm{\delta\mu}_{n}.$ (3.11e)
The results converge towards the solution preferred by the data, independently from the choice of the initial condition [41, 42, 39, 43]. Therefore, starting from different initial conditions, the results of the iterative smoothing method will approach this solution by different paths. We collect all reconstructions with a better $\chi^{2}$ than the best-fit ${\Lambda\text{CDM}}$, which then make up a non-exhaustive set of solutions.
@@
where denotes $\mathrm{d}/\mathrm{d}z$. The formalism of the smooth derivative is described in the appendix.
@@
It is worth noting that, while we can obtain $\mathcal{D}^{\prime}$ and $\mathcal{D}^{\prime\prime}$ from the smoothing of the supernovae directly, to obtain $h$ and $h^{\prime}$, one needs to know $\Omega_{k,0}$. Therefore, wrongly assuming flatness will imply an error when inferring $h(z)$, $h^{\prime}(z)$, and $q(z)$. This is just a restatement of the $\mathcal{O}_{k}$ test.
Figure 4: The bottom rows show our estimations of $H_{0}r_{\text{d}}/c$ for our two methods at the BAO data points for the three fiducial cosmologies. The central line is the collection of central values of $H_{0}r_{\text{d}}$ for each reconstruction, and the coloured band is the one-sigma error from the BAO around the central value. The top rows show likelihood for estimate (3.9b) at fixed iteration, and the third row for estimate (3.9a). The colour bars of the second and middle rows are the same as one of the corresponding smooth figures (1 to 2).
Figure 5: Curvature parameter: $\Theta$. The bottom row shows $\Theta=h\mathcal{D}^{\prime}$ for our three fiducial cosmologies. Like in Fig. 4, the central value is obtained by combining the BAO with each iteration of the smooth SN reconstructions, and the error-band from the $1\sigma$ errors from the BAO. The top row shows the likelihood of $\Theta$ for each smooth reconstruction. We note that in the middle panel of the top row, there is no true value for $\Theta$ since $\Theta(z)=\sqrt{1+\Omega_{k}\mathcal{D}^{2}}$ varies with redshift.
@@ 4.1 Supernova Smoothing
Fig. 1 shows the results of the smoothing for the flat-${\Lambda\text{CDM}}$ model. Each line is an iteration of the smoothing color-coded by its $\Delta\chi^{2}$ with respect to the best-fit ${\Lambda\text{CDM}}$ model. The top-left panel shows the reconstructed distance moduli with the mock SNIa data. The middle- and bottom-left panels show the comoving distance $\mathcal{D}(z)/\mathcal{D}_{\text{fid}}(z)$ and the expansion history (${}_{\text{fid}}(z)/hh(z)$ normalized to the fiducial values. The middle- and bottom-right panels show $O_{m}$ and the deceleration parameter $q$. We are able to reconstruct the fiducial quantities.
@@ 4.2 $H_{0}r_{\text{d}}$
The bottom row of Fig. 4 shows our two estimations of $H_{0}r_{\text{d}}/c$ for the each model from eqs (3.9a) and  (3.9b). At each BAO data point, we plot the central value and the ±$\pm 1\sigma$ band for each reconstruction with $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\Lambda\text{CDM}}<2.3$. It is worth stressing here that at every BAO point, we have $N$ estimations of $H_{0}r_{\text{d}}/c$, corresponding to $N$ reconstructions of the distance-redshift relations from the SNIa (combined with the mock BAO data point). Each of these estimates also has an associated error bar, shown as an error band on the figure. At each redshift point, the reconstructed $H_{0}r_{\text{d}}/c$ are re-sorted by increasing central value for aesthetic reasons. The fiducial value $H_{0}r_{\text{d}}^{\text{fid}}/c=0.03435$ is shown as a dashed line. For the three fiducial cosmologies, the estimations from eq. (3.9a) are consistent with $H_{0}r_{\text{d}}^{\text{fid}}$.
The top and middle rows show the likelihood of $\Theta$ and $\mathcal{O}_{k}$ for a given iteration for eq. (3.9b) and  (3.9a). This is equivalent to calculating the weighted average over the different redshift bins weighted by $1/\sigma_{i}^{2}$. The weighted average of a quantity $\bm{A}=(A(z_{1}),\dots,A(z_{n}))$ with covariance matrix $C_{A}$ over the different redshift bins and its variance are given by:
$\langle A\rangle$ $=\frac{\mathbbm{1}^{\top}\bm{C_{A}}^{-1}\bm{A}}{\mathbbm{1}^{\top}\bm{C_{A}}^{-1}\mathbbm{1}},$ (4.1a)
$\sigma^{2}_{\langle A\rangle}$ $=\left(\mathbbm{1}^{\top}\bm{C_{A}}^{-1}\mathbbm{1}\right)^{-1}.$ (4.1b)
@@
In the case of $A=H_{0}r_{\text{d}}$, the correlation between two different redshift bins is negligible, and the covariance matrix in eq. (4.1) is diagonal. In the case of eq. 3.9a (middle row), the three fiducial cosmologies yield similar reconstructions, regardless of the curvature and DE. It is worth paying attention to the $\Omega_{k,0}=0.1$ case (middle columns). While estimator (3.9a) (third row) is totally consistent with the fiducial value, eq. (3.9b) is centred around 0.035, and most of the reconstructions are over $2\sigma$ away from the fiducial value, showing an inconsistency between the two estimators. This comes from the fact that we do not calculate $1/h$ but $\mathcal{D}^{\prime}$, which do not coincide in a non-flat FLRW universe, as seen above. Indeed, we can see in the first row that the values of $\mathcal{D}^{\prime}Hr_{\text{d}}/c$ diverge from the fiducial values at high redshift, reflecting equation (3.18). The minimal and maximal values of the precision of the weighted average are reported in table 2, with typically 0.5% precision.
Quantity flat-${\Lambda\text{CDM}}$ $\Omega_{k,0}=0.1$ PEDE
$\langle\mathcal{D}r_{\text{d}}/d_{\text{C}}\rangle$ 0.52% 0.52% 0.52%
$\langle\mathcal{D}^{\prime}Hr_{\text{d}}/c\rangle$ 0.51% 0.51% 0.51%
$\mathcal{O}_{k}$ 0.032–0.034 0.034–0.037 0.030–0.033
$\Theta$ 0.0074–0.0075 0.0075–0.0077 0.0074–0.0076
Table 2: Precision for our three fiducial cosmologies. The precision for our two estimates of $H_{0}r_{\text{d}}$ is relative to the central value, while the precisions in $\mathcal{O}_{k}$ and $\Theta$ are absolute. For $\mathcal{O}_{k}$ and $\Theta$, we quote the minimum and maximum values. For $H_{0}r_{\text{d}}$, the difference between the minimum and maximum values of the precision is negligible.
@@ 4.3 Curvature Test
The Bottom panel of Fig. 5 shows our estimate of $\Theta(z)=h\mathcal{D}^{\prime}=\sqrt{1+\Omega_{k,0}^{2}\mathcal{D}}$ for the three cosmologies. The top panel shows the weighted average of $\Theta(z)$ over $z$ for a given iteration of the smoothing, colour-coded by $\Delta\chi^{2}$ of the SNIa reconstruction. For a flat FLRW universe, we have $\Theta(z)\equiv 1$. Indeed, for the Flat-$\Lambda$ (left) and PEDE (right) cases, our estimates are consistent with $\Theta=1$. The interesting case is the $\Omega_{k,0}=0.1$ case, where $\Theta$ is not constant, and we can see a trend of $\Theta(z)$ deviating from 1 as $z$ increases. Similarly, the weighted average (peak of the likelihood) is larger than in the other two cases, and only marginally consistent with $\Theta=1$.
Fig. 6 shows similar plots for the $\mathcal{O}_{k}$ parameter. Similarly to $\Theta$, the Flat-$\Lambda$ and PEDE case are perfeclty consistent with flatness ($\Omega_{k,0}=0$), while the $\Omega_{k,0}=0.1$ case is only marginally consistent with $\Omega_{k,0}=0$ and perfectly consistent with $\Omega_{k,0}=0.1$, showing the constraining power of this test.
@@ 5 Summary and Conclusion
We combined forecast BAO data for a DESI 5-year-like survey and a realistic forecast for LSST SNIa to assess the ability of the joint surveys to provide model-independent litmus tests of the flat ${\Lambda\text{CDM}}$ model. For this purpose, we generated mock data for three cosmologies: flat ${\Lambda\text{CDM}}$  a non-flat ${\Lambda\text{CDM}}$ universe with $\Omega_{k,0}=0.1$, and an example of dynamical dark energy, the PEDE model.
@@
The expected quality of LSST SNIa will allow to reconstruct very accurately certain diagnostics such as $Om$ and $q$. In addition, we quantify the ability of the joint surveys to constrain $H_{0}r_{d}/c$. We then proceed to testing the FLRW metric, measuring $\Theta$ and $O_{k}$. The combination of LSST and DESI 5-years will allow to constrain $O_{k}$ to within $\sim 0.035$, and $\Theta$ to within 0.0075. It is worth noting that the credible interval reported by DESI DR1 is $(-0.078,0.068)$ for the flat ${\Lambda\text{CDM}}$ model without any external data. We show that with the expected five-year data accuracy, a model-independent precision on the curvature parameter will be competitive.
In the case of the $\Omega_{k,0}=0.1$ universe, the flatness assumption of eq. (3.3) is now incorrect, which leads to a biased expansion history and inconsistent estimations of $H_{0}r_{\text{d}}/c$ from our two methods. Therefore, $H_{0}r_{\text{d}}$ itself can be seen as a curvature test. An inconsistency between the two estimates of $H_{0}r_{\text{d}}$, or with $\mathcal{O}_{k}(z)=1$ or $\Theta(z)=1$ can just be interpreted either as evidence of departure from flatness, or inconsistency between the two data sets.

### v3

@@ 4.2 Iterative Smoothing
The smoothing width is set to $\Delta=0.3$ following previous analyses in [48, 49, 31, 50, 32, 51, 52, 53]. $W(z)$ can be understood as a Gaussian smoothing kernel of the variable $u=\ln(1+z)$, and the denominator as a normalization factor.
We define the $\chi^{2}$ value of the reconstruction $\hat{\mu}_{n}(z)$ at iteration $n$ as
$\chi^{2}_{\text{SN},n}=\bm{\delta\mu}_{n}^{\intercal}\cdot\bm{C}_{\text{SN}}^{-1}\cdot\bm{\delta\mu}_{n}.$ (4.7)
The results converge towards the solution preferred by the data, independently from the choice of the initial condition [48, 49, 31, 32]. Therefore, starting from different initial conditions, the results of the iterative smoothing method will approach this solution by different paths. More precisely, we start from the best-fit $k$-${\Lambda\text{CDM}}$, flat ${\Lambda\text{CDM}}$, and flat ${\Lambda\text{CDM}}$ with fixed $\Omega_{\text{m},0}=\{0,0.1,\dots,1\}$. We collect all reconstructions with $\Delta\chi^{2}_{\text{SN},n}=\chi^{2}_{\text{SN},n}-\chi^{2}_{\text{SN,LCDM}}<2.3$ relative to the best-fit ${\Lambda\text{CDM}}$. These reconstructions then make up a non-exhaustive set of plausible distances and expansion histories.
@@
where denotes $\mathrm{d}/\mathrm{d}z$. The derivatives of $\mu$ are directly obtained during the smoothing process of the supernovae. The formalism of the smooth derivative is described in the appendix.
@@
It is worth noting that, while we can obtain $\mathcal{D}^{\prime}$ and $\mathcal{D}^{\prime\prime}$ from the smoothing of the supernovae directly, to obtain $h$ and $h^{\prime}$, one needs to know $\Omega_{k,0}$. Therefore, wrongly assuming flatness will imply an error when inferring $h(z)$, $h^{\prime}(z)$, and $q(z)$. This is just a restatement of the $\mathcal{O}_{k}$ test. In addition, it is worth noticing that, while BAO provide both separate (but correlated) measurement of $d_{\text{M}}/r_{\text{d}}$ and $d_{\text{H}}/r_{\text{d}}$, $\mathcal{D}$, $\mathcal{D}^{\prime}$ and $\mathcal{D}^{\prime\prime}$ are obtained from the same measurements.
Figure 5: The bottom row shows our estimations of $c/H_{0}r_{\text{d}}$ for our two methods (Eqs. (4.4a) and (4.4b)) at the BAO data points for the four fiducial cosmologies. The black dotted line is the fiducial value. The central line is the collection of central values of $c/H_{0}r_{\text{d}}$ for each reconstruction, and the coloured band is the one-sigma error from the BAO around the central value. The top rows show the likelihood for estimate (4.4a) at fixed iteration, and the second row for estimate (4.4b), and the third row shows the combination of both, taking into account the correlation. The colour bars of the top three rows are the same as that of the corresponding smooth figures (1 to 4).
Figure 6: Curvature parameter: $\Theta$. The bottom row shows $\Theta=h\mathcal{D}^{\prime}$ for our four fiducial cosmologies. Like in Fig. 5, the central value is obtained by combining the BAO with each iteration of the smooth SN reconstructions, and the error-band from the $1\sigma$ errors from the BAO. The top row shows the likelihood of $\Theta$ for each smooth reconstruction. We note that in the second and fourth panels of the top row, there is no true value for $\Theta$ since $\Theta(z)=\sqrt{1+\Omega_{k,0}\mathcal{D}^{2}}$ varies with redshift.
@@ 5.1 Supernova Smoothing
Fig. 1 shows the results of the smoothing for the flat-${\Lambda\text{CDM}}$ fiducial model. Each line is an iteration of the smoothing color-coded by its $\Delta\chi^{2}$ with respect to the best-fit ${\Lambda\text{CDM}}$ model. The top-left panel shows the reconstructed distance moduli with the mock SNIa data. The middle- and bottom-left panels show the comoving distance $\mathcal{D}(z)/\mathcal{D}_{\text{fid}}(z)$ and the inverse of the expansion history $h_{\text{fid}}(z)/h(z)$ normalized to the fiducial values. Again, the colored lines come from the smoothing of the SNIa, while the data points with error-bars come from the BAO. The middle- and bottom-right panels show $\mathcal{O}m$ and the deceleration parameter $q$. We are able to reconstruct the fiducial quantities. We report the $\Delta\chi^{2}_{\text{fid}}=\chi^{2}_{\text{fid}}-\chi^{2}_{{\Lambda\text{CDM}}}$ in the last column of Table 1.
@@ 5.2 $H_{0}r_{\text{d}}$
The bottom row of Fig. 5 shows our two estimations of $c/H_{0}r_{\text{d}}$ from eqs (4.4a) and  (4.4b) for each fiducial model. At each BAO data point, we plot the central value and the ±$\pm 1\sigma$ band for each reconstruction $n$ with $\Delta\chi_{\text{SN},n}^{2}=\chi_{\text{SN},n}^{2}-\chi^{2}_{\text{SN, ${\Lambda\text{CDM}}$}}<2.3$. It is worth stressing here that at every BAO point, we have $N_{\text{rec}}$ estimations of $c/H_{0}r_{\text{d}}$, corresponding to $N_{\text{rec}}$ reconstructions of the distance-redshift relations from the SNIa (combined with the mock BAO data point). Each of these estimates also has an associated error bar, shown as an error band on the figure. At each redshift point, the reconstructed $c/H_{0}r_{\text{d}}$ are re-sorted by increasing central value for aesthetic reasons. The fiducial value $c/H_{0}r_{\text{d}}^{\text{fid}}=29.73$ is shown as a dashed line. For the four fiducial cosmologies, the estimations from eq. (4.4a) are consistent with $H_{0}r_{\text{d}}^{\text{fid}}$.
Rows 1 to 3 show the (normalized) likelihood for iteration $n$
$\mathcal{L}_{n,i}(x)$ $\propto\exp\left(-\frac{\chi^{2}_{n}}{2}\right)\mbox{, where}$ (5.1a)
$\chi^{2}_{n}$ $=-\frac{1}{2}(x\mathbbm{1}-\bm{X}_{i})^{\intercal}\cdot\bm{C}_{i}^{-1}\cdot(x\mathbbm{1}-\bm{X}_{i}),$ (5.1b)
$\bm{X}_{1}^{\intercal}$ $=\begin{pmatrix}\frac{1}{\mathcal{D}(z_{1})}\frac{d_{\text{M}}(z_{1})}{r_{\text{d}}},\dots,\frac{1}{\mathcal{D}(z_{N})}\frac{d_{\text{M}}(z_{N})}{r_{\text{d}}}\end{pmatrix},$ (5.1c)
$\bm{X}_{2}^{\intercal}$ $=\begin{pmatrix}\frac{1}{\mathcal{D}^{\prime}(z_{1})}\frac{d_{\text{H}}(z_{1})}{r_{\text{d}}},\dots,\frac{1}{\mathcal{D}^{\prime}(z_{N})}\frac{d_{\text{H}}(z_{N})}{r_{\text{d}}}\end{pmatrix},$ (5.1d)
$\bm{X}_{3}$ $=\begin{pmatrix}\bm{X}_{1}\\ \bm{X}_{2}\end{pmatrix}$ (5.1e)
@@
are the data vectors, and $N=11$ is the number of BAO data points. The error propagation, including the covariance between the two BAO modes, is detailed in Appendix B. In the case of eq. (4.4a) (row 1), the three fiducial cosmologies yield similar reconstructions, regardless of the curvature and DE. This is to be expected, since both distances from the SNIa and BAO should agree regardless of the curvature. It is worth paying attention to the $k$-$\Lambda$ and $k$-PEDE cases (second and fourth columns). While estimator (4.4a) (row 1) is totally consistent with the fiducial value, eq. (4.4b) (row 2) is centred around 28 to 29, and most of the reconstructions are over $2\sigma$ away from the fiducial value, showing an inconsistency between the two estimators. This comes from the fact that while the radial mode of the BAO provides us with $1/h$, the SNIa provide us with $\mathcal{D}^{\prime}$, which do not coincide in a non-flat FLRW universe, as seen above. Indeed, we can see in the bottom row that the values of $d_{\text{H}}/(r_{\text{d}}\mathcal{D}^{\prime})$ diverge from the fiducial values at high redshift, reflecting equation (4.14). For each quantity $A$, we report the minimal and maximal values of the precision $\sigma_{A}$ and the spread $\Delta A=(A-A_{\text{fid}})$ taken over all reconstructions in Table 2. Namely, for a given reconstruction $n$, we obtain the likelihood in the top three panels, and obtain the corresponding precision and spread. We then report the minimal and maximal value over the whole reconstructions. In the case of $X_{i}$ (eqs. (5.1c) to (5.1e), we normalize the result by the fiducial value. The third row shows a greater precision by combining the two estimators despite the correlation between the radial and transverse BAO modes.
Quantity flat-$\Lambda$ $k$-$\Lambda$ PEDE $k$-PEDE
$\sigma_{X_{1}}/X_{\text{fid}}$ 0.52% 0.52% 0.52 – 0.53% 0.52%
$\Delta{X_{1}}/X_{\mathrm{fid}}$0.44% – 0.44% -0.77% – 0.68% -0.05% – 0.98% -0.37% – 0.49%
$\sigma_{X_{2}}/X_{\text{fid}}$ 0.50 – 0.52% 0.49 – 0.51% 0.50 – 0.52% 0.49 – 0.51%
$\Delta{X_{2}}/X_{\mathrm{fid}}$1.55% – 1.94% -4.40% – 0.72% -1.74% – 2.04% -3.82% – -0.18%
$\sigma_{X_{3}}/X_{\mathrm{fid}}$ 0.28 – 0.29 % 0.28% 0.28 – 0.29% 0.28%
$\Delta{X_{3}}/X_{\mathrm{fid}}$0.68% – 1.06% -1.96% – 0.47% -0.75% – 1.24% -1.82% – 0.02%
$\sigma_{\mathcal{O}_{k}}$ 0.037 – 0.040 0.040 – 0.045 0.035 – 0.038 0.039 – 0.042
$\Delta{\mathcal{O}_{k}}$ 0.004 – 0.015 -0.095 – -0.031 0.004 – 0.016 -0.095 – -0.055
$\sigma_{\Theta}$ 0.009 0.009 0.009 0.009
$\Delta\Theta$0.026 – 0.029 -0.0277 – 0.0332
Table 2: Table of precision and spread for our four fiducial cosmologies. The reported values are the minimum and the maximum values of the $1\sigma$ CL of the likelihoods in Figs. 5 to 7 over all reconstructions. The spread of quantity $A$, noted as $\Delta A$ is defined as the difference between the peak of the likelihood and the fiducial value. For simplicity, we note $X=c/H_{0}r_{\text{d}}$, and the indices refer to the different estimators (see eqs. (5.1c) to (5.1e)). The precisions for our three estimates of $X=c/H_{0}r_{\text{d}}$ are relative to the true value, while the precisions and spreads in $\mathcal{O}_{k}$ and $\Theta$ are absolute. We note that in the non-flat cases ($k$-$\Lambda$ and $k$-PEDE), there is no fiducial value for $\Theta$ since $\Theta$ evolves with redshift, so we do not report the spread.
@@ 5.3 Curvature Test
The bottom panel of Fig. 6 shows our estimate of $\Theta(z)=h\mathcal{D}^{\prime}=\sqrt{1+\Omega_{k,0}\mathcal{D}^{2}}$ for the three cosmologies. The top panel shows the normalized likelihood of $\Theta(z)$ over $z$ for a given iteration of the smoothing, colour-coded by $\Delta\chi^{2}$ of the SNIa reconstruction. As expected, for the Flat-$\Lambda$ (left) and PEDE (right) cases, our estimates are consistent with $\Theta=1$. The interesting cases are the $k$-$\Lambda$ and $k$-PEDE cases (second and fourth columns). We can see in the bottom panel a trend of $\Theta(z)$ deviating from 1 as $z$ increases. Similarly, in the top panel, the likelihoods peak at higher values than in the other two cases, and only marginally consistent with $\Theta=1$.
Fig. 7 shows similar plots for the $\mathcal{O}_{k}$ parameter. Similarly to $\Theta$, the Flat-$\Lambda$ and PEDE case are perfectly consistent with flatness ($\Omega_{k,0}=0$), while the $\Omega_{k,0}=0.1$ case is only marginally consistent with $\Omega_{k,0}=0$ and perfectly consistent with $\Omega_{k,0}=0.1$, showing the constraining power of this test.
@@
In order to assess the power and significance of the test, we perform a hypothesis test. One must bear in mind that the problem is not symmetric here: the null hypothesis $H_{0}$ to test is “$\Omega_{k0}=0$”. For the two flat models, we are interested in type I errors: what is $p_{I}$, the probability of wrongly rejecting flatness? For the two non-flat models, the question is “What is the probability of failing to reject flatness?”, i.e., type II errors, with probability $p_{II}$.
To assess these, we resimulate 1000 realizations of the data. Since different iterations of the smoothing are correlated, one cannot simply combine them. Instead, we choose for each realization the one with the lowest $\Delta\chi^{2}$. We then calculate, for each realization, the likelihood of $\Omega_{k,0}\pm\sigma_{\Omega_{k,0}}$. For a given value of the significance $\alpha$, we want to measure $p_{I}$ and $p_{II}$, assuming a Gaussian likelihood for $\Omega_{k,0}$, i.e., $p_{I}=2P(X>|\Omega_{k,0}|)$, and $p_{II}=1-2P(X>|\Omega_{k,0}|)$. The results are reported on Table 2. While the probabilities of type I errors are reasonable, for a given $\alpha$, the probability of type II errors are typically 40% ($\alpha=0.05$) to 60% ($\alpha=0.01$). While a LSST 3 year+DESI 5 years survey is not sufficient to rule out flatness at a high confidence level for curvature values of 0.1, it is important to keep in mind that this is only using low-redshift data, and does not need any assumption regarding the dark energy model.
$\Lambda$ $k$-$\Lambda$ PEDE $k$-PEDE
$\alpha$ $p_{I}$ $p_{II}$ $p_{I}$ $p_{II}$
0.05 0.068 0.363 0.065 0.334
0.01 0.016 0.595 0.014 0.548
Table 3: Probabilities of type I and type II errors for the four fiducial models.
@@ 6 Summary and Conclusion
We combined forecast BAO data for a DESI 5-year-like survey and a realistic forecast for LSST SNIa to assess the ability of the joint surveys to provide model-independent litmus tests of the flat ${\Lambda\text{CDM}}$ model. For this purpose, we generated mock data for four cosmologies: flat ${\Lambda\text{CDM}}$  a non-flat ${\Lambda\text{CDM}}$ universe with $\Omega_{k,0}=0.1$, and two examples of dynamical dark energy, the PEDE model, with $\Omega_{k,0}=0$ and $\Omega_{k,0}=0.1$.
@@
The expected quality of LSST SNIa will allow to reconstruct very accurately certain diagnostics such as $\mathcal{O}m$ and $q$. In addition, we quantify the ability of the joint surveys to constrain $c/H_{0}r_{\text{d}}$. We then proceed to testing the FLRW metric, measuring $\Theta$ and $\mathcal{O}_{k}$. The combination of LSST 3-year and DESI 5-years will allow to constrain $O_{k}$ to within $\sim 0.035$, and $\Theta$ to within 0.0075, with a spread due to the SNIa uncertainties of up to ±$\pm 0.1$ and 0.03 respectively. It is worth noting that the credible interval on $\Omega_{k,0}$ reported by DESI DR1 is $(-0.078,0.068)$ for the flat ${\Lambda\text{CDM}}$ model without any external data. We show that with the expected five-year data accuracy, a model-independent precision on the curvature parameter will be competitive.
In the case of the $\Omega_{k,0}=0.1$ universes, the flatness assumption of eq. (2.3) is now incorrect, which leads to a biased expansion history and inconsistent estimations of $c/H_{0}r_{\text{d}}$ from our two methods. Therefore, $H_{0}r_{\text{d}}$ itself can be seen as a curvature test. An inconsistency between the two estimates of $H_{0}r_{\text{d}}$, or with $\mathcal{O}_{k}(z)=$ constant or $\Theta(z)=1$ can just be interpreted either as evidence of departure from flatness, or inconsistency between the two data sets.